OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (lgbasallote):

use point slope form \[\huge y - y_1 = m (x - x_1)\] where m is the slope x1 is the x coordinate y1 is the y oordinate does that help?

OpenStudy (anonymous):

would it look like this ? 5=-1(1)

OpenStudy (lgbasallote):

close.. y + 5 = -1(x-1) dont omit y and x they are VERY important

OpenStudy (lgbasallote):

now simplify that

OpenStudy (anonymous):

y+5=-x+1 ?

OpenStudy (anonymous):

x+y=-4 ?

OpenStudy (anonymous):

thankyou i got it!

OpenStudy (lgbasallote):

glad you got it